OpenStudy (anonymous):

The area of a given hexagon is equal to the area of an equilateral triangle whose perimeter is 36 inches. Find the length of a side of the regular hexagon.

OpenStudy (anonymous):

 Here, the 2 similar equilateral triangle have Area scale factor 6, so linear scale factor root(6), so the length is 12*(1/root6)=2root6. I may be wrong.

OpenStudy (anonymous):

Solve for x in: \[ \frac{\sqrt{3}} 4 \times (12)^2 = \frac{\sqrt{3}} 4 \times x^2 \times 6 \]

OpenStudy (anonymous):

\(\implies x= 2\sqrt{6}\)

OpenStudy (anonymous):

So, @henpen is right.

OpenStudy (anonymous):

Area of the equilateral triangle of perimeter 36 is \( \frac{\sqrt{3}} 4 \times (12)^2 \)

OpenStudy (anonymous):

Simmilary area of an equilateral triangle is given by \(\frac{\sqrt{3}} 4 \times x^2 \times 6\)

OpenStudy (anonymous):

where x is the side length.